OpenStudy (anonymous):

Solve using elimination: -7y + 2x - 8 = 0 -4x + 14y = 14 Answer Infinitely many solutions No solution (-3, -2) (11, 2)

OpenStudy (anonymous):

Change the first equation to line up the x,y, and constant with the second. You should end up with the equation 2x-7y=8 -4x+14y=14 Now, multiply the top equation by 2. 4x-14y=16 -4x+14y=14 Eliminate your x -14y=16 14y=14 Eliminate your y 0=30

OpenStudy (loser66):